OpenStudy (anonymous):

Simplify. Assume that no denominator is 0 and that 0^0 is not considered. 10n^7 / 15n^3

OpenStudy (anonymous):

\[\frac{10n^7}{15n^3}\] \[10n^7 = 2 \times 5 \times n \times n \times n \times n \times n \times n \times n\] \[15n^3 = 3 \times 5 \times n \times n \times n\] can you tell me what in between those two have in common?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

and how many n?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

yup so they are both divisible by 5n^3, this is the largest common factor \[\huge \frac{10n^7 \div 5n^3}{15n^3 \div 5n^3}\] what will you get?

OpenStudy (anonymous):

2n^10/ 3n^6 ?

OpenStudy (anonymous):

you subtract when dividing exponents so it would be 2n^4/3

OpenStudy (anonymous):

wow woops my bad didn't see the notification -_- \[\huge n^m \div n^n = n^{m-n}\] so yea that would leave you with 2n^4/3

OpenStudy (anonymous):

thank you